# TWISTED CONJUGACY IN DIHEDRAL ARTIN GROUPS I: TORUS KNOT GROUPS

### v2

@@ INTRODUCTION
For a finitely generated group $G$ with generating set $X$, we say two elements $u,v\in G$ are twisted conjugate, by some automorphism $\phi\in\mathrm{Aut}(G)$, if there exists an element $w\in G$ such that $v=\phi^{-1}(w)uw$. The twisted conjugacy problem (TCP) asks whether there exists an algorithm to determine if two elements, given as words over $X$, are twisted conjugate by some automorphism $\phi\in\mathrm{Aut}(G)$. Whilst the twisted conjugacy problem is a relatively new decision problem, in comparison to the conjugacy problem, several important results have been established (see [1, 3, 6, 9]). Note that a positive solution to the twisted conjugacy problem implies a positive solution to the conjugacy problem, however the converse is not necessarily true [2, Corollary 4.9].
In a series of two papers, we solve the TCP for dihedral Artin groups. Artin groups are defined by a simple graph with edge labelling from $\mathbb{Z}_{\geq 2}$. Despite their long history of study, there are still open questions, such as the decidability of the word problem, which are not known for all Artin groups. A typical approach to understand these open problems is to consider subclasses of Artin groups, which are often defined using conditions on the defining graph. Dihedral Artin groups are one such class, and are defined by a simple graph with precisely two vertices and one edge, that is, they have a group presentation $G(m)=\langle a,b\mid{}_{m}(a,b)={}_{m}(b,a)\rangle$, where $m\geq 3$ and (${}_{m}(a,b)$ is the word $abab\dots$ of length $m$.
These one-relator, torsion free groups are appealing to study from a computational perspective, and recent studies have found interesting results such as decidability of equations [5], and solvable conjugacy problem in linear time [10]. We note here that the twisted conjugacy problem for dihedral Artin groups was already known to be solvable by Juhász [11], with respect to length preserving automorphisms. To solve the twisted conjugacy problem for dihedral Artin groups with respect to all automorphisms, we consider when the edge labelling of our defining graph is either odd or even. In this first paper, we focus on odd dihedral Artin groups and further improve on Juhász’s result, by constructing an algorithm to check if two words, representing group elements, are twisted conjugate. Moreover, this algorithm has linear time complexity, based on the length of the input words.
@@
thmTCP The $\mathrm{TCP}(G(m))$, where $m$ is odd, $m\geq 3$, is solvable in linear time. The key step is to utilise the fact that any odd dihedral Artin group $G(m)$ is isomorphic to a torus knot group. In particular, $G(m)$ is isomorphic to the group with presentation $\langle x,y\mid x^{2}=y^{m}\rangle$, where $m$ is odd. With this presentation, we find that the automorphism group of $G(m)$ is finite of order two. Moreover, we can use a geodesic normal form derived by Fujii [7], to construct a linear time algorithm to solve the $\mathrm{TCP}(G(m))$.
As an application of Section 1, we consider the conjugacy problem in extensions of odd dihedral Artin groups, using a criteria from [2] (see 4.1). This criteria requires understanding orbit decidability for subgroups $A\leq\mathrm{Aut}(G)$, for some group $G$. The orbit decidability problem asks whether we can determine if for two elements $u,v\in G$, there exists an automorphism $\phi\in A$ such that $v$ is conjugate to $\phi(u)$. To apply the criteria from [2], we require what is known as the ‘action subgroup’ of odd dihedral Artin groups to be orbit decidable. For odd dihedral Artin groups, we prove a stronger statement, by showing that all subgroups of the automorphism group of odd dihedral Artin groups are orbit decidable.
@@
thmextension Let $G=G(m)\rtimes H$ be an extension of an odd dihedral Artin group by a finitely generated group $H$ which satisfies conditions $(ii)$ and $(iii)$ from 4.1 (e.g. let $H$ be torsion-free hyperbolic). Then $G$ has decidable conjugacy problem. We note that the conjugacy problem is solvable for virtual dihedral Artin groups [5, Corollary 3.2], which proves Section 1 in the case where $H$ is finite.
The structure of this paper is as follows. After providing necessary details on decision problems, twisted conjugacy and odd dihedral Artin groups in Section 2, we construct a linear time solution to the TCP for odd dihedral Artin groups in Section 3. Finally in Section 4, we study orbit decidability in odd dihedral Artin groups, which allows us to use the criteria from [2] to find new examples of group extensions of odd dihedral Artin groups with solvable conjugacy problem.
@@
All groups in this paper are finitely generated, and all finite generating sets are inverse-closed.
@@
    The (uniform) twisted conjugacy problem for $G$, denoted $\mathrm{TCP}(G)$, takes as input two words $u,v\in X^{\ast}$, and decides whether $u$ and $v$ represent groups elements which are $\phi$-twisted conjugate in $G$, for some $\phi\in\mathrm{Aut}(G)$.
Clearly a solution to the $\mathrm{TCP}(G)$ implies a solution to $\mathrm{TCP}_{\phi}(G)$ for all $\phi\in\mathrm{Aut}(G)$, and therefore a solution to the $\mathrm{CP}(G)$ and the $\mathrm{WP}(G)$. We first observe that to solve the $\mathrm{TCP}(G)$, we can study the outer automorphism group, rather than the automorphism group as a whole. This was originally noted in [9].
@@
Let $\psi\circ\phi\in\mathrm{Aut}(G)$, where $\psi\in\mathrm{Inn}(G)$, $\phi\in\mathrm{Out}(G)$. If $\phi$ is trivial, then the $\mathrm{TCP}_{\psi}(G)$ is equivalent to solving the $\mathrm{CP}(G)$ by the following relations:
@@
for some $g\in G$. Otherwise, we have
@@
for some $g\in G$. Therefore if $\phi\in\mathrm{Out}(G)$ is non-trivial, then the $\mathrm{TCP}_{\psi\circ\phi}(G)$ is equivalent to solving the $\mathrm{TCP}_{\phi}(G)$, for $\phi\in\mathrm{Out}(G)$.
@@
For any automorphism $\phi\in\mathrm{Aut}(G)$ of a group $G$, we say $\phi$ is:
@@
    length preserving if $l(w)=l(\phi(w))$ for any word $w\in X^{\ast}$, and
  2. (ii)
@@
[10, Proposition 3.1] The conjugacy problem is solvable in dihedral Artin groups in linear time.
@@
Let $G(m)$ denote the dihedral Artin group, with presentation given in Eq. 1. For $m$ odd,
@@
by setting $x={}_{m}(a,b),y=ab$. Here $G(m)$ is isomorphic to a torus knot group.
For notation we let $X=\{x,y\}$, and $m=2k+1$ for some $k\in\mathbb{Z}_{\geq 1}$. Before considering twisted conjugacy, we first need to understand the outer automorphisms of this group.
@@
[8, Theorem C] Let $G(m)$ be defined as Eq. 2. Then $\mathrm{Out}(G(m))\cong C_{2}$.
@@
ALGORITHM FOR THE TWISTED CONJUGACY PROBLEM
To solve the twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$, where $\phi\in\mathrm{Out}(G(m))$ as in Eq. 3, we will show that every geodesic is twisted conjugate to a minimal length twisted conjugacy representative which is Garside free. Moreover, for two minimal representatives which are twisted conjugate, we will show that they are related by a finite sequence of $\phi$-cyclic permutations and equivalent geodesics. Since $\phi$ is length-preserving, this will lead to our basic algorithm, which we later improve to give a linear time algorithm.
@@
    For $c\in\{-1,0,1\}$, $x^{\epsilon_{1}}vx^{\epsilon_{2}}\Delta^{c}\rightarrow\begin{cases}v\Delta^{c},&\epsilon_{1}=\epsilon_{2},\\ v\Delta^{c+\epsilon},&\epsilon_{1}\neq\epsilon_{2}.\end{cases}$
@@
    For $c\in\{-1,0,1\}$, $y^{b_{1}}vy^{b_{n}}\Delta^{c}\rightarrow\begin{cases}y^{b_{1}-b_{n}}v\Delta^{c}&b_{1}-b_{n}\in[-(m-1),m-1]\cap\mathbb{Z}_{\neq 0},\\ v\Delta^{c},&b_{1}=b_{n},\\ v\Delta^{c+\epsilon},&|b_{1}-b_{n}|=m.\end{cases}$
@@
[7, Proposition 3.5] Let $u\in(3)$ be geodesic. Then the following conditions must all be satisfied:
@@
The $\mathrm{TCP}_{\phi}(G(m))$, where $m$ is odd, $m\geq 3$, with respect to $\phi\colon x\mapsto x^{-1},y\mapsto y^{-1}$, is solvable.
@@
Step 1: Geodesic form {adjustwidth}1.5cm Write $u,v$ in geodesic forms $\overline{u},\overline{v}$ respectively, given by Table 1. Step 2: Simple cases {adjustwidth}1.5cm If $\overline{u}$ or $\overline{v}$ is of the form $x^{a}$ or $y^{b}$, then apply Lemma 3.3 to determine if $\overline{u}$ and $\overline{v}$ are twisted conjugate. Otherwise, we can assume $\overline{u},\overline{v}$ contain at least one $x$ and $y$ factor. Step 3: Twisted cyclic geodesics {adjustwidth}1.5cm Apply $\phi$-cyclic permutations and geodesic reductions to obtain $\overline{u}_{\min},\overline{v}_{\min}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, such that $\overline{u}\sim_{\phi}\overline{u}_{\min}$ and $\overline{v}\sim_{\phi}\overline{v}_{\min}$ (existence follows from 3.2 and Lemma 3.10). If $l(\overline{u}_{\min})\neq l(\overline{v}_{\min})$, then Output = False (by 3.15). Step 4: Set of representatives {adjustwidth}1.5cm Compute all possible sequences of $\phi$-cyclic permutations and equivalent words from $\overline{u}_{\min}$, to get a finite set $\mathcal{D}$ of minimal length twisted conjugacy representatives. If $\overline{v}_{\min}\in\mathcal{D}$, then Output = True. Otherwise, Output = False (by 3.15). ∎
@@
Let $\overline{u}_{\min}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, where $l(\overline{u}_{\min})=n$. Then the number of minimal length group elements $v\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, such that $u\sim_{\phi}v$, is bounded above by $2n(n+1)$.
@@
where each ##\## represents an $x^{\pm 1}$ term. Note if $u$ starts with a $y$ factor, then we can apply a $\phi$-cyclic permutation to obtain the correct form. Suppose the number of $x$ terms in $u$ equals $q\leq m$. Consider the following $\phi$-cyclic permutation:
@@
Here the number of $x$ terms in $u^{\prime}$ must be equal to either $q-1$ or $q+1$, depending on whether we move an $x$ or $x^{-1}$ from the last block in $u$. In particular, each time we apply a $\phi$-cyclic permutation of a block, we either increase or decrease the number of $x$ terms by 1. Moreover, the number of $x$ terms in any element is bounded below by 0 and above by $m$. By repeated application of $\phi$-cyclic permutations of blocks, we see that for any $v$ obtained from $\overline{u}_{\min}$ via $\phi$-cyclic permutations must have either an even or odd number of $x$ terms.
@@
In total, we obtain $2(m+1)\sum^{m}_{i}|b_{i}|$ group elements, by computing all $\phi$-cyclic permutations of $\overline{u}_{\min}$ using this block decomposition. To see this, we group together block decomposition’s which only differ by one $\phi$-cyclic permutation of a block. These pairs have either odd or even number of $x$ terms, which gives us $m+1$ group elements. Summing across all $y$ powers, to include all possible $\phi$-cyclic permutations, we get $(m+1)\sum^{m}_{i}|b_{i}|$ group elements. This is then doubled when we repeat this procedure for $\overline{u}_{\min}$, with respect to a block decomposition of the form $u=\left(y^{b_{1}}\#\right)\dots\left(y^{b_{m}}\#\right)$, which gives us a total of $2(m+1)\sum^{m}_{i}|b_{i}|$ group elements. Since $m<n$, this value has an upper bound of $2n(n+1)$. ∎
@@
This already improves our algorithm to quadratic time, in that at Step 4, we can compute all minimal length group elements which are twisted conjugate to $\overline{u}_{\min}$, and then check if $\overline{v}_{\min}$ represents one of these group elements or not. This can be improved further to obtain a linear time Step 4 of our algorithm as follows.
@@
Step 4: Checking group elements {adjustwidth}1.5cm For $u=x^{a_{1}}y^{b_{1}}\dots x^{a_{\tau}}y^{b_{\tau}}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, define
$X(u):=\begin{cases}(0,2,4,\dots,\tau)&\#x\text{-terms in}\;u\;\text{is even},\\ (1,3,5,\dots,\tau-1)&\#x\text{-terms in}\;u\;\text{is odd},\end{cases}$
@@
if $\tau$ is even, and define analogously for $m$ odd. We compute all $\phi$-cyclic permutations of $\overline{u}_{\min}\mathsf{CycGeo}_{\phi,\overline{(3)}}$ using the block decomposition, whilst keeping track of how many $x$ terms can exist for each element we obtain (for each $\phi$-cyclic permutation, we switch between odd and even). For each group element $w\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ obtained via $\phi$-cyclic permutations, we check whether the exponents of $y$ factors of $w$ and $\overline{v}_{\min}$ match exactly, and that $X(w)=X(\overline{v}_{\min})$. If this holds for some $w$, then Output = True. Otherwise, if no such $w$ exists, then Output = False . We are now able to determine linear complexity for our adapted algorithm.

### v3

@@ INTRODUCTION
For a finitely generated group $G$ with inverse-closed generating set $X$, we say two elements $u,v\in G$ are twisted conjugate, by some automorphism $\phi\in\mathrm{Aut}(G)$, if there exists an element $w\in G$ such that $v=\phi^{-1}(w)uw$. The twisted conjugacy problem (TCP) asks whether there exists an algorithm to determine if two elements, given as words over $X$, are twisted conjugate by some automorphism $\phi\in\mathrm{Aut}(G)$. Whilst the twisted conjugacy problem is a relatively new decision problem, in comparison to the conjugacy problem, several important results have been established (see [2, 6, 9, 12]). Note that a positive solution to the twisted conjugacy problem implies a positive solution to the conjugacy problem, however the converse is not necessarily true [3, Corollary 4.9].
In a series of two papers, we solve the twisted conjugacy problem for dihedral Artin groups. Artin groups are defined by a simple graph with edge labelling from $\mathbb{Z}_{\geq 2}$. Despite their long history of study, there are still open questions, such as the decidability of the word problem, which are not known for all Artin groups. A typical approach to understand these open problems is to consider subclasses of Artin groups, which are often defined using conditions on the defining graph. Dihedral Artin groups are one such class, and are defined by a simple graph with precisely two vertices and one edge, that is, they have a group presentation $G(m)=\langle a,b\mid{}_{m}(a,b)={}_{m}(b,a)\rangle$, where $m\geq 3$ and (${}_{m}(a,b)$ is the word $abab\dots$ of length $m$.
These one-relator, torsion free groups are appealing to study from a computational perspective, and recent studies have found interesting results such as decidability of equations [8], and solvable conjugacy problem in linear time [13]. We note here that the twisted conjugacy problem for dihedral Artin groups was already known to be solvable by Juhász [14], with respect to length-preserving automorphisms. To solve the twisted conjugacy problem for dihedral Artin groups with respect to all automorphisms, we consider when the edge labelling of our defining graph is either odd or even. In this first paper, we focus on odd dihedral Artin groups and further improve on Juhász’s result, by constructing an algorithm to check if two words, representing group elements, are twisted conjugate. Moreover, this algorithm has linear time complexity, based on the length of the input words.
@@
thmTCP The twisted conjugacy problem $\mathrm{TCP}(G(m))$ for dihedral Artin groups, where $m$ is odd, $m\geq 3$, is solvable in linear time. The key step is to utilise the fact that any odd dihedral Artin group $G(m)$ is isomorphic to a torus knot group. In particular, $G(m)$ is isomorphic to the group with presentation $\langle x,y\mid x^{2}=y^{m}\rangle$, where $m$ is odd. With this presentation, we find that the automorphism group of $G(m)$ is finite of order two. Moreover, we can use a geodesic normal form derived by Fujii [10], to construct a linear time algorithm to solve the $\mathrm{TCP}(G(m))$.
As an application of Section 1, we consider the conjugacy problem in extensions of odd dihedral Artin groups, using a criteria from [3] (see 5.1). This criteria requires understanding orbit decidability for subgroups $A\leq\mathrm{Aut}(G)$, for some group $G$. The orbit decidability problem asks whether we can determine if for two elements $u,v\in G$, there exists an automorphism $\phi\in A$ such that $v$ is conjugate to $\phi(u)$. To apply the criteria from [3], we require what is known as the ‘action subgroup’ of odd dihedral Artin groups to be orbit decidable. For odd dihedral Artin groups, we prove a stronger statement, by showing that all subgroups of the automorphism group of odd dihedral Artin groups are orbit decidable.
@@
thmextension Let $G=G(m)\rtimes H$ be an extension of an odd dihedral Artin group by a finitely generated group $H$ which satisfies conditions $(ii)$ and $(iii)$ from 5.1 (e.g. let $H$ be torsion-free hyperbolic). Then $G$ has decidable conjugacy problem. We note that the conjugacy problem is solvable for virtual dihedral Artin groups [8, Corollary 3.2], which proves Section 1 in the case where $H$ is finite.
The structure of this paper is as follows. After providing necessary details on decision problems, twisted conjugacy and odd dihedral Artin groups in Section 2, we construct a linear time solution with respect to outer automorphisms in Section 3. This is then extended in Section 4 to solve the TCP for odd dihedral Artin groups. Finally in Section 5, we study orbit decidability, which allows us to use the criteria from [3] to find new examples of group extensions of odd dihedral Artin groups with solvable conjugacy problem.
@@
All groups in this paper are finitely generated, and all finite generating sets are inverse-closed. For a subset $S$ of a group, we let $S^{\pm}=S\cup S^{-1}$, where $S^{-1}=\{s^{-1}\mid s\in S\}$. In certain cases, group presentations will be stated using generating sets which are not inverse-closed, and when we pass to the relevant inverse-closed generating set, we won’t explicitly mention the corresponding change in the group presentation.
@@
    The (uniform) twisted conjugacy problem for $G$, denoted $\mathrm{TCP}(G)$, takes as input two words $u,v\in X^{\ast}$ and $\phi\in\mathrm{Aut}(G)$, and decides whether $u$ and $v$ represent groups elements which are $\phi$-twisted conjugate in $G$.
A solution to the $\mathrm{TCP}(G)$ implies a solution to $\mathrm{TCP}_{\phi}(G)$ for all $\phi\in\mathrm{Aut}(G)$, and therefore a solution to the $\mathrm{CP}(G)$ and the $\mathrm{WP}(G)$. We first observe that to solve the $\mathrm{TCP}(G)$, we can study the outer automorphism group, rather than the automorphism group as a whole. This was originally noted in [12].
@@
Let $\psi\circ\phi\in\mathrm{Aut}(G)$, where $\psi\in\mathrm{Inn}(G)$, $\phi\in\mathrm{Out}(G)$ are fixed. If $\phi$ is trivial, then the $\mathrm{TCP}_{\psi}(G)$ is equivalent to solving the $\mathrm{CP}(G)$ by the following relations:
@@
for some fixed $g\in G$. Otherwise, we have
@@
for some fixed $g\in G$. Therefore if $\phi\in\mathrm{Out}(G)$ is non-trivial, then the $\mathrm{TCP}_{\psi\circ\phi}(G)$ is equivalent to solving the $\mathrm{TCP}_{\phi}(G)$, for some fixed $\phi\in\mathrm{Out}(G)$.
@@
Let $G=\langle X\rangle$ be a group. We say $\phi\in\mathrm{Aut}(G)$ is:
@@
    length-preserving if $|\pi(w)|=|\phi(\pi(w))|$ for any word $w\in X^{\ast}$, and
  2. (i)
@@
[13, Proposition 3.1] The conjugacy problem is solvable in dihedral Artin groups in linear time.
@@
Let $G(m)$ denote the dihedral Artin group, with presentation given in Eq. 1. For $m=2k+1$ for some $k\in\mathbb{Z}_{\geq 1}$,
@@
by setting $x={}_{m}(a,b),y=ab$. Here $G(m)$ is isomorphic to a torus knot group. For notation we let $X=\{x,y\}^{\pm}$, which we refer to as the free product generating set.
Before considering twisted conjugacy, we first need to understand the outer automorphisms of this group.
@@
[11, Theorem C] Let $G(m)$ be defined as in Eq. 2. Then $\mathrm{Out}(G(m))\cong C_{2}$.
@@
ALGORITHM FOR OUTER AUTOMORPHISMS
Our main goal is to solve the twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$, where $\phi\in\mathrm{Out}(G(m))$ as in Eq. 3. We will then generalise this algorithm, using 2.3, to solve the twisted conjugacy problem $\mathrm{TCP}(G(m))$.
We will first show that when $\phi\in\mathrm{Out}(G(m))$ as in Eq. 3, then every geodesic is twisted conjugate to a minimal length twisted conjugacy representative which is Garside free. Moreover, for two minimal representatives which are twisted conjugate, we will show that they are related by a finite sequence of $\phi$-cyclic permutations and equivalent geodesics. Since $\phi$ is length-preserving, this will lead to our basic algorithm, which we later improve to give a linear time algorithm.
@@
    For $c\in\{0,1\}$, $x^{\epsilon_{1}}vx^{\epsilon_{2}}\Delta^{c}\rightarrow\begin{cases}v\Delta^{c},&\epsilon_{1}=\epsilon_{2},\\ v\Delta^{c+\epsilon},&\epsilon_{1}\neq\epsilon_{2}.\end{cases}$
@@
    For $c\in\{0,1\}$, $y^{b_{1}}vy^{b_{n}}\Delta^{c}\rightarrow\begin{cases}y^{b_{1}-b_{n}}v\Delta^{c}&b_{1}-b_{n}\in[-(m-1),m-1]\cap\mathbb{Z}_{\neq 0},\\ v\Delta^{c},&b_{1}=b_{n},\\ v\Delta^{c+\epsilon},&|b_{1}-b_{n}|=m.\end{cases}$
@@
[10, Proposition 3.5] Let $u\in(3)$ be geodesic. Then the following conditions must all be satisfied:
@@
The twisted conjugacy problem $\mathrm{TCP}_{\phi}(G(m))$, where $m$ is odd, $m\geq 3$, with respect to $\phi\colon x\mapsto x^{-1},y\mapsto y^{-1}$, is solvable.
@@
Step 1: Geodesic form {adjustwidth}1.5cm Write $u,v$ in geodesic forms $\overline{u},\overline{v}$ respectively, given by Table 2. Step 2: Simple cases {adjustwidth}1.5cm If $\overline{u}$ or $\overline{v}$ is of the form $x^{a}$ or $y^{b}$, then apply Lemma 3.3 and 3.15 to determine if $\overline{u}$ and $\overline{v}$ are twisted conjugate. In particular, $\overline{u}\sim_{\phi}\overline{v}$ if and only if:
$\begin{cases}\overline{u}=x^{a},\;\overline{v}=x^{b},&a,b\;\text{have the sameparity},\\ \overline{u}=y^{t+mc},\;\overline{v}=y^{s+md},&\{c,t\},\{d,s\}\;\text{both have opposite parity or same parity},\\ \overline{u}=x^{a},\;\overline{v}=y^{t+mc},&a\;\text{even},\{c,t\}\;\text{havesame parity}.\end{cases}$
Otherwise, we can assume $\overline{u},\overline{v}$ contain at least one $x$ and $y$ factor. Step 3: Twisted cyclic geodesics {adjustwidth}1.5cm Apply $\phi$-cyclic permutations and geodesic reductions to obtain $\overline{u}_{\min},\overline{v}_{\min}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, such that $\overline{u}\sim_{\phi}\overline{u}_{\min}$ and $\overline{v}\sim_{\phi}\overline{v}_{\min}$ (existence follows from 3.2 and Lemma 3.10). If $l(\overline{u}_{\min})\neq l(\overline{v}_{\min})$, then Output = False (by 3.15). Step 4: Set of representatives {adjustwidth}1.5cm Compute all possible sequences of $\phi$-cyclic permutations and equivalent words from $\overline{u}_{\min}$, to get a finite set $\mathcal{D}$ of minimal length twisted conjugacy representatives. If $\overline{v}_{\min}\in\mathcal{D}$, then Output = True. Otherwise, Output = False (by 3.15). ∎
@@
Let $\overline{u}_{\min}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, where $l(\overline{u}_{\min})=n$. Then the number of minimal length group elements $v\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, such that $u\sim_{\phi}v$, is bounded above by $n(n+1)$.
@@
where each ##\## represents an $x^{\pm 1}$ term. Note if $u$ starts with a $y$ factor, then we can apply a $\phi$-cyclic permutation to obtain the correct form. Suppose the number of $x$ terms in $u$ equals $q\leq m$. Consider the following $\phi$-cyclic permutation of $u$:
@@
Here the number of $x$ terms in $u^{\prime}$ must be equal to either $q-1$ or $q+1$, depending on whether we move an $x$ or $x^{-1}$ from the last block in $u$. In particular, each time we apply a $\phi$-cyclic permutation of a block, we either increase or decrease the number of $x$ terms by 1. Moreover, the number of $x$ terms in any element is bounded below by 0 and above by $m$.
@@
In total, we obtain $(m+1)(m+\sum^{m}_{i}|b_{i}|)$ group elements, by computing all $\phi$-cyclic permutations of $\overline{u}_{\min}$ using this block decomposition. To see this, we group together block decomposition’s which only differ by one $\phi$-cyclic permutation of a block. These pairs have either odd or even number of $x$ terms, which gives us $m+1$ group elements. Summing across all $y$ powers, to include all possible $\phi$-cyclic permutations, we get $(m+1)(m+\sum^{m}_{i}|b_{i}|)$ group elements. Since $m<n$, this value has an upper bound of $n(n+1)$. ∎
@@
This already improves our algorithm to $\mathcal{O}(n^{3})$, in that at Step 4, we can compute all minimal length group elements which are twisted conjugate to $\overline{u}_{\min}$, and then check if $\overline{v}_{\min}$ represents one of these group elements or not. This can be improved further to obtain a linear time Step 4 of our algorithm as follows.
@@
Step 4: Checking group elements {adjustwidth}1.5cm For $w=x^{a_{1}}y^{b_{1}}\dots x^{a_{\tau}}y^{b_{\tau}}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, define
$X(w):=\begin{cases}(0,2,4,\dots,\tau)&\#x\text{-terms in}\;u\;\text{is even},\\ (1,3,5,\dots,\tau-1)&\#x\text{-terms in}\;u\;\text{is odd},\end{cases}$
@@
if $\tau$ is even, and define analogously for $m$ odd. We compute all $\phi$-cyclic permutations of $\overline{u}_{\min}\mathsf{CycGeo}_{\phi,\overline{(3)}}$ using the block decomposition, whilst keeping track of how many $x$ terms can exist for each element we obtain (for each $\phi$-cyclic permutation, we switch between odd and even). For each group element $w\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ obtained via $\phi$-cyclic permutations from $\overline{u}_{\min}$, we check whether the exponents of $y$ factors of $w$ and $\overline{v}_{\min}$ match exactly, and that $X(w)=X(\overline{v}_{\min})$. If this holds for some $w$, then Output = True. Otherwise, if no such $w$ exists, then Output = False .
Example 3.20.
Recall 3.17. Here $m=2$ and $(m+1)(m+\sum^{m}_{i=1}|b_{i}|)=3(2+2)=12$ from 3.19, and so there are 12 minimal length group elements within the set $\mathcal{D}$. We list them in Table 1, by taking the minimal shortlex representative for each group element (where $x<x^{-1}<y<y^{-1})$, and also recording the parity of the number of $x$-terms in each group element.
Group element Parity of $x$ terms Group element Parity of $x$ terms
$xy^{-1}x^{-1}y$ Odd $y^{-1}xy^{-1}x^{-1}$ Odd
$xy^{-1}xy^{-1}$ Even $x^{-1}y^{-1}x^{-1}y^{-1}$ Even
$yxy^{-1}x$ Even $yx^{-1}y^{-1}x^{-1}$ Even
$xyx^{-1}y^{-1}$ Odd $yxyx^{-1}$ Odd
$x^{-1}yx^{-1}y$ Even $xyxy$ Even
$y^{-1}x^{-1}yx^{-1}$ Even $y^{-1}xyx$ Even
Table 1: Group elements from 3.17
We apply our adapted Step 4 to the words $\overline{u}_{\min}=xy^{-1}x^{-1}y$, $\overline{v}_{\min}=yx^{-1}y^{-1}x^{-1}$. Here $\overline{u}_{\min}\sim_{\phi}\overline{v}_{\min}$ (see Fig. 1). After computing all possible $\phi$-cyclic permutations from $\overline{u}_{\min}$, we obtain 8 words:
$xy^{-1}x^{-1}y\leftrightarrow y^{-1}xy^{-1}x^{-1}\leftrightarrow xy^{-1}xy^{-1}\leftrightarrow yxy^{-1}x$
$\leftrightarrow x^{-1}yxy^{-1}\leftrightarrow yx^{-1}yx\leftrightarrow x^{-1}yx^{-1}y\leftrightarrow y^{-1}x^{-1}yx^{-1}.$
Consider $w=yxy^{-1}x$ from this list. Then the $y$-exponents of $w$ and $\overline{v}_{\min}$ match exactly, and $X(w)=X(\overline{v}_{\min})=\{0,2\}$, and so our algorithm will output True as required. Alternatively, consider $\overline{v}_{\min}=yxy^{-1}x^{-1}$. Again by Fig. 1, $\overline{u}_{\min}\not\sim_{\phi}\overline{v}_{\min}$. Here $\{0,2\}=X(w)\neq X(\overline{v}_{\min})=\{1\}$, and since $w$ is the only word from this list with matching $y$-exponents to $\overline{v}_{\min}$, then our algorithm will output False as required.
We are now able to determine linear complexity for our adapted algorithm.